OpenStudy (anonymous):

an undiscovered planet, many lightyears from Earth, has one moon in a periodic orbit. this moon takes 1960*10^3 s(about 23 days) on average to complete one nearly circular revolution around the unnamed planet. if the distance from the center of the moon to the surface of the planet is 215.0*10^6m and the planet has a radius of 3.40*10^6m, calculate the moon's radial acceleration. i've used 2(pi)/t...is that right??

OpenStudy (anonymous):

2pi/T will give you the angular velocity of the moon

OpenStudy (anonymous):

so do i use a = v²/R?

OpenStudy (anonymous):

a+v^2/R?

OpenStudy (anonymous):

*=...not +

OpenStudy (anonymous):

well if you've already worked out the angular velocity, call it omega, why not use the formula r x (omega squared) for the radial acceleration

OpenStudy (anonymous):

v squared over r is correct, but you may as well work with the omega formula

OpenStudy (anonymous):

\[a _{r}=v^2/r=r \omega^2\] either way is fine

OpenStudy (anonymous):

do i have to find the total distance first? i did that and got= 2.184*10^8 then i did 2(pi)/t t=1960*10^3 and got =3.21/s i then entered that into a=Rw^2 and got 2250.4*10^6 as the final answer....am i doing it correct?

OpenStudy (anonymous):

first, yes you are right, the radius we need is the sum of the planet radius plus the surface to moon distance you're calculation of omega is not quite correct - you are missing an exponent

OpenStudy (anonymous):

\[\frac{ 2\pi }{ 1960e10^3 }=?\]

OpenStudy (anonymous):

.0644...right?

OpenStudy (anonymous):

erm, no 2pi/1960000

OpenStudy (anonymous):

oh so it'll be 3.206e^-6?

OpenStudy (anonymous):

correct now you just need to calculate r times omega squared for the radial acceleration

OpenStudy (anonymous):

so i should get 55.64*10^-5?

OpenStudy (anonymous):

what do you have for r ?

OpenStudy (anonymous):

the total distance...

OpenStudy (anonymous):

yes, the distance from centre of planet to centre of moon

OpenStudy (anonymous):

yes it was 218.4*10^6....thats supposed to be r right?

OpenStudy (anonymous):

that is right, yes and what do you have for omega squared ?

OpenStudy (anonymous):

3.06e^-6

OpenStudy (anonymous):

that is omega

OpenStudy (anonymous):

yes..so squared it would be 10.28e^-6?

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

what is 10^-6 squared ?

OpenStudy (anonymous):

you can't just square the first few numbers and forget about the power of ten, that has to be squared as well !

OpenStudy (anonymous):

oh ok, so it would be 1.97e^-5

OpenStudy (anonymous):

what would be ?

OpenStudy (anonymous):

omega

OpenStudy (anonymous):

omega we already calculated to be 3.2e10-6

OpenStudy (anonymous):

now we need omega squared

OpenStudy (anonymous):

r= 218.4*10^6 and omega squared =1.97e^-5

OpenStudy (anonymous):

???????????????????????

OpenStudy (anonymous):

omega square = omega multiplied by omega if omega is 3.2e10-6, just work out omega squared for me

OpenStudy (anonymous):

i get 1.024e^13

OpenStudy (anonymous):

you need practise doing simple calculations with numbers in scientific notation if omega = 3.2e10-6 then omega squared = 1.024e10-11

OpenStudy (anonymous):

you have to take great care to be accurate doing these numerical calculations

OpenStudy (anonymous):

where are you getting e10-11 from?

OpenStudy (anonymous):

i'm just entering it into my calculator

OpenStudy (anonymous):

the power of 10 in omega is minus 6 if you square it, that will give you a power of minus 12, but 3.2 squared is about 10, so instead of writing 10 times 10 to power -12, you write 1 times 10 to power -11

OpenStudy (anonymous):

oooh ok, i see now

OpenStudy (anonymous):

it's very important to be able to do that kind of calculation correctly, otherwise you might understand the physics but not be able to get the correct numbers out

OpenStudy (anonymous):

so now you have r and you have omega squared - can you calculate the radial accereration for me ?

OpenStudy (anonymous):

i have 82273146.83

OpenStudy (anonymous):

oh my god, where did that come from

OpenStudy (anonymous):

r = 218.4 x 10^6 and omega squared = 1.02 x 10^-11, just multiply those two numbers together

OpenStudy (anonymous):

i hit the wrong button, sorry..the correct answer is .00222768

OpenStudy (anonymous):

that's more like it : ) so the radial acceleration is just over 2mm per second squared.

OpenStudy (anonymous):

yes..thanks so much, seems i'm making careless mistakes in my calculations

OpenStudy (anonymous):

it's easily done - a good night's sleep can help

OpenStudy (anonymous):

yes that's true i haven't slept in couple of days now, thanks again

OpenStudy (anonymous):

welcome